y=-4-4x
I need to find the Y intercept and then I need to find the X intercept. I am not nowing how to work the problem out the right way.

Found 3 solutions by rfer, scott8148, stanbon:
Answer by rfer(16322)

x int is when y=0
y int is when x=0
y=-4-4x
0=-4-4x
4=-4x
x=-1
(-1,0)
--------------
y=-4-4*0
y=-4
(0,-4)

Answer by scott8148(6628)

at the y-intercept, x equals zero ___ and at the x-intercept, y equals zero

y-int ___ y = -4 -4(0) ___ y = -4

x-int ___ (0) = -4 - 4x ___ 4 = -4x ___ -1 = x

Answer by stanbon(75887)

y=-4-4x
I need to find the Y intercept and then I need to find the X intercept. I am not knowing how to work the problem out the right way.
---------------
To find the y-intercept, let x = 0; then solve for "y":
y = -4-4x
Let x = 0
y = -4-4*0
y = -4-0
y = -4
The y-intercept is the point (0,-4)
==============================
To find the x-intercept, let y = 0; then solve for "x":
y = -4-4x
Let y = 0
-4-4x = 0
Add 4x to both sides to get:
-4 = 4x
Divide both sides by 4 to get:
x = -1
---
The x-intercept is the point (-1,0)
====================================
Cheers,
Stan H.
===============